### v2

# Reconstruction-Free Anomaly Detection with Diffusion Models

Abstract
Despite the remarkable success, recent reconstruction-based anomaly detection (AD) methods via diffusion modeling still involve fine-grained noise-strength tuning and computationally expensive multi-step denoising, leading to a fundamental tension between fidelity and efficiency. In this paper, we propose a novel inversion-based AD approach – “detection via noising in latent space” – which circumvents explicit reconstruction. Importantly, we contend that the limitations in prior reconstruction-based methods originate from the prevailing “detection via denoising in RGB space” paradigm. To address this, we model AD under a reconstruction-free formulation, which directly infers the final latent variable corresponding to the input image via DDIM inversion, and then measures the deviation based on the known prior distribution for anomaly scoring. Specifically, in approximating the original probability flow ODE using the Euler method, we only enforce very few inversion steps to noise the clean image to pursue inference efficiency. As the added noise is adaptively derived with the learned diffusion model, the original features for the clean testing image can still be leveraged to yield high detection accuracy. We perform extensive experiments and detailed analysis across three widely used image AD datasets under the unsupervised unified setting to demonstrate the effectiveness of our model, regarding state-of-the-art AD performance, and about $2\times$ inference time speedup without diffusion distillation.
Projecthttps://SkyShunsuke.github.io/InversionADProject/
Codehttps://github.com/SkyShunsuke/InversionAD
Introduction
Image anomaly detection (AD) focuses on identifying images that deviate from normal patterns, with critical applications in industrial defect detection and automated medical diagnosis (J et al. 2020; Bercea et al. 2023). In real-world scenarios, anomalous samples are often scarce and costly to acquire. Consequently, researchers began focusing on unsupervised AD, where the detection model was trained solely on normal images to identify potential anomalies (Roth et al. 2021; Batzner, Heckler, and König 2023).
Due to the robust capabilities for approximating distributions, diffusion models (Ho, Jain, and Abbeel 2020; Sohl-Dickstein et al. 2015) have emerged as powerful tools for AD by precisely modeling normal data distribution (Zhang et al. 2023b; Yao et al. 2024; Lu et al. 2023a; Zhang et al. 2023a; Mousakhan, Brox, and Tayyub 2023; He et al. 2024; Wyatt et al. 2022; Fučka, Zavrtanik, and Skočaj 2025; Teng et al. 2023; Feng et al. 2025; Beizaee et al. 2025). The prevailing paradigm in these methods is to treat anomalies as noise: by adding noise of a certain intensity to an anomalous image and then denoising it with a diffusion model trained solely on normal data, the anomalous regions can be effectively reconstructed as normal regions. The anomaly score is therefore computed by measuring the mean squared error (MSE) between the original input and its reconstructed counterpart, as shown in Figure 1 (a).
We argue that the above “detection via denoising in RGB space” paradigm is less effective given the following fundamental limitations: (1) Noise-strength sensitivity. The detection performance depends heavily on the selection of appropriate noise strengths. Stronger noise corrupts normal regions, increasing false positives, while weaker noise leads to overly well-recovered anomaly regions (Lu et al. 2023a; Yao et al. 2024); (2) Computational expense in multi-step denoising. Satisfactory reconstruction via diffusion models often requires iterative denoising steps, which is significantly time-consuming for real-world applications. As summarized in Table 1, most existing reconstruction-based methods involve high inference latency.
(a) Conventional reconstruction-based AD paradigm
(b) Our inversion-based AD paradigm
Figure 1: Conceptual comparison of conventional and our proposed AD paradigm. and Conventional reconstruction-based paradigm (a) first perturbs an input sample $\mathbf{x}_{0}$ to a latent state $\mathbf{x}_{t}$ at step $t$ , and then denoises $\mathbf{x}_{t}$ back to $\mathbf{x}_{0}$ . The anomaly score is computed as the mean squared error (MSE) between the original input and its reconstructed sample. In contrast, our inversion-based paradigm (b) directly infers the latent state at the final step, $\mathbf{x}_{T}$, by tracing the PF-ODE trajectories. The anomaly score is then determined based on the typicality of $\mathbf{x}_{T}$ within the tractable latent distribution.
Therefore, we introduce the “detection via noising in latent space” paradigm, which bypasses the reconstruction process in prior AD approaches. Our core insight is that, since the learned diffusion model is solely aware of the normal data, we can utilize such learned normal data distribution through latent inversion, rather than reconstruction, as shown in Figure 1 (b). Specifically, for a given image, we directly infer the corresponding latent variable at the final diffusion step by tracing the probability flow ordinary differential equation (PF-ODE (Song et al. 2021)) trajectory. Since PF-ODE is by nature deterministic, this constructs a one-to-one correspondence between the noise prior distribution and the distribution of normal images, ensuring that normal images can be ideally mapped to focus on a high-density region. Conversely, anomalous images will be mapped to low-density regions. Therefore, we can detect anomalies by evaluating the degree of deviation of the inferred latent variable under the prior distribution.
While the above method detours noise strength tuning, it does not straightforwardly lead to a reduction of detection speed. In particular, the high inference latency primarily stems from the high-order ODE solver introduced for accurate reconstruction from inverted latents. To address this, we are inspired by (Teng et al. 2023; Heng, Thiery, and Soh 2024) to measure the norm of the score function as a proxy for typicality, which is robust to the step size selection of the Euler method. This allows us to enforce very few inversion steps to diffuse the clean image to pursue inference efficiency. Moreover, it contributes to a significantly more efficient inference paradigm compared to prior approaches. Importantly, because the learned diffusion model enables noise to be adaptively added during inversion, the key features of the original clean testing image can still be leveraged to harvest high detection accuracy. As our method is designed for the inference stage, it can be applied in a plug-and-play fashion to be incorporated into existing AD models. Experimental evaluations demonstrate that our method substantially improves inference latency, yet still yields high AD performance, across four widely-used AD benchmarks.
Our core contributions include three folds:
  • Identifying the inherent limitations of existing reconstruction-based methods in achieving an optimal efficiency-accuracy balance, which we attribute fundamentally to the “detection via denoising in RGB space” paradigm.
  • Proposing a novel paradigm InvAD, detection via noising in latent space, with an inversion-based AD framework that eliminates reconstruction requirements by directly estimating latent variables from inputs and quantifying anomaly scores through deviation measurements under the prior distribution.
  • Demonstrating that the inversion-based approach, which infers latent variables, is simpler, faster, and achieves superior detection performance compared to conventional reconstruction-based methods.
Background
Table 1: Comparison of the properties of diffusion-based AD methods. Normal-only means whether the method involves pseudo-anomalies in training, NFE stands for the number of function evaluations, and TS refers to the timestep.
Method Normal-only NFE FPS TS Tuning-free Multi-class Scoring scheme
DiffAD [ICCV’23] N/A N/A Mask prediction
DiAD [AAAI’24] 1010 1.5 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
GLAD [ECCV’24] 750750 0.2 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
TransFusion [ECCV’24] 2020 1.6 Mask prediction
DAD [TPAMI’25] $\{2,400\}$ 9.1 Mask prediction
DDAD [GCPR’24] $\{5,10,25\}$ 0.8 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
ScoreDD [ArXiv’22] 1515 N/A $\nabla_{\mathbf{x}}p_{\theta}(\mathbf{x}_{t})||_{2}$
OmiAD [ICML’25] 11 39.4 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
DeCo-Diff [CVPR’25] 1010 17.0 $\text{MSE}(g_{\phi}(\mathbf{x}_{0}),g_{\phi}(\hat{\mathbf{x}}_{0}))$
InvAD (Ours) 33 88.1 $\log p_{\theta}(\mathbf{x}_{T})$
Diffusion Models
Diffusion modeling (Ho, Jain, and Abbeel 2020; Song et al. 2021) is a recently emerged generative modeling framework, which aims to reverse a Markov perturbation process (forward process). In general, each transition of the forward process is defined as:
$q\bigl{(}\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\bigr{)}$ $:=\mathcal{N}\!\Bigl{(}\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\,\mathbf{x}_{t-1},\bigl{(}1-\frac{\alpha_{t}}{\alpha_{t-1}}\bigr{)}\mathbf{I}\Bigr{)},$ (1)
where monotonically decreasing sequence $\alpha_{1:T}\in(0,1]^{T}$ schedules the noise. The sampling procedure can then be achieved with a reverse process, which defines a Markov process with a Gaussian transition kernel:
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}):=\mathcal{N}\bigl{(}\mathbf{x}_{t-1};\bm{\mu}_{\theta}^{(t)}(\mathbf{x}_{t}),\bigl{(}1-\frac{\alpha_{t}}{\alpha_{t-1}}\bigr{)}\mathbf{I}\bigr{)},$ (2)
where $\theta$ denotes the learnable parameter of the DDPM. The training is realized by minimizing the $\epsilon$-parameterized objective:
$\mathcal{L}_{\gamma}\left(\epsilon_{\theta}\right)\coloneqq\sum_{t=1}^{T}\gamma_{t}\mathbb{E}_{\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right),\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})-\epsilon_{t}\right\|_{2}^{2}\right],$ (3)
where $\mathbf{x}_{t}$ is sampled from $q_{t}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$, and $\mathbf{\gamma}\coloneqq[\gamma_{1},\ldots,\gamma_{T}]^{T}$ denotes the weighting vector. During sampling, one can first sample a pure noise from $p_{\theta}(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$, and then progressively trace Equation 2 for denoising.
Denoising Diffusion Implicit Models
The Gaussian approximation in Equation 2 generally works satisfactorily when the total number of diffusion steps, $T$, is sufficiently large (Ho, Jain, and Abbeel 2020). However, increasing $T$ also magnifies the number of function evaluations (NFE), which slows down the sampling procedure. To address this, Song et al. introduced denoising diffusion implicit models (DDIM), allowing for faster sampling by leveraging a subset of the original diffusion process (Song, Meng, and Ermon 2021). Specifically, they reformulate the reverse process as follows:
$\begin{aligned} \mathbf{x}_{\tau_{i\!-\!1}}&=\sqrt{\alpha_{\tau_{i\!-\!1}}}f_{\theta}(\mathbf{x}_{\tau_{i}})+\sqrt{1-\alpha_{\tau_{i\!-\!1}}}\,\epsilon_{\theta}^{(\tau_{i})}\bigl{(}\mathbf{x}_{\tau_{i}}\bigr{)},\end{aligned}$
where $f_{\theta}(\mathbf{x}_{\tau_{i}})=\mathbf{x}_{\tau_{i}}-\sqrt{1-\alpha_{\tau_{i}}}\,\epsilon_{\theta}^{(\tau_{i})}\bigl{(}\mathbf{x}_{\tau_{i}}\bigr{)}/\sqrt{\alpha_{\tau_{i}}}$ and $\tau_{S}=[\tau_{1},\tau_{2},\ldots,\tau_{S}=T]\subset\{1,2,\ldots,T\}$. This deterministic sampling equation can be interpreted as the Euler integration of the corresponding PF-ODE:
$\mathrm{d}\mathbf{y}_{t}=\epsilon_{\theta}^{(t)}\,\mathrm{d}p_{t},$ (5)
where $\mathbf{y}_{t}\coloneqq\mathbf{x}_{t}/\sqrt{\alpha_{t}}$ and $p_{t}\coloneqq\sqrt{1/{\alpha_{t}}-1}$. The PF-ODE gradually transforms a noise distribution into a data distribution, and vice versa. This approach, known as inversion, maps a given data sample back to its noise representation.
Diffusion-Based Anomaly Detection
The above diffusion-based generative models have been recently employed to handle AD. Existing diffusion-based AD methods generally follow the reconstruction-based paradigm. In particular, to address the challenge of noise level selection inherent in this framework, these approaches can be broadly classified into the following three categories:
  • Conditional information injection-based methods incorporate auxiliary information, such as class labels (He et al. 2024), semantic features (Mousakhan, Brox, and Tayyub 2023; He et al. 2024), or unmasked regions (Feng et al. 2025; Beizaee et al. 2025; Xu, Wang, and Du 2024). As pointed out in (Mousakhan, Brox, and Tayyub 2023), conditioning the denosing process with such additional information contributes to reducing reconstruction error in the normal region and increasing it for anomalous regions. However, since this line of methods conditions the diffusion model at training time, they can lead to increased training complexity.
  • Discriminator-based methods introduce an additional discriminator, typically trained to distinguish between normal and anomalous pixels (Zhang et al. 2023b, a). While this line of methods enables efficient anomaly localization, they rely on pseudo-anomalies, biasing AD towards specific types of anomalies.
  • Adaptive timesteps-based methods attempt to adjust or identify perturbation timesteps by comparing the reconstruction errors of denoised results across different time steps (Yao et al. 2024). As the above two types of methods do not incorporate an explicit mechanism for adjusting the perturbation time step, it is typically treated as a hyperparameter or addressed via ensembling over multiple time steps (Lu et al. 2023a). Details can be found in the ”timestep tuning” column of Table 1.
In contrast to the state-of-the-art non-diffusion method (Batzner, Heckler, and König 2023) that achieves over 200 FPS, current diffusion-based methods mostly reach 1 FPS on average. While a recent diffusion-based approach (Feng et al. 2025) tackles this issue by introducing score distillation to constrain the diffusion iteration to only one step, the training for distillation can be complicated due to the instability of the adversarial learning. Table 1 summarizes the properties of current diffusion-based AD methods. Overall, a simple, efficient, yet effective paradigm for diffusion-based AD remains unexplored to date.
Methodology
To address the limitations of current diffusion-based methods, we introduce a novel diffusion-based AD framework with latent inversion, where the sampling trajectory is reformulated as $\mathbf{x}_{0}\rightarrow\mathbf{x}_{T}$. In contrast to conventional reconstruction-based approaches $\bigl{(}\mathbf{x}_{0}\rightarrow\mathbf{x}_{t}\rightarrow\hat{\mathbf{x}}_{0}\bigr{)}$, our method infers the final latent state $\mathbf{x}_{T}$ directly from $\mathbf{x}_{0}$ by following the learned PF-ODE trajectory. Figure 1(b) illustrates our reconstruction-free AD paradigm, which leverages the inversion process to achieve efficient yet accurate AD.
Inversion-Based Anomaly Detection
Given an image $\mathbf{x}$ and a score function $\epsilon_{\theta}^{(t)}$ trained using Equation 3, our goal is to infer the latent representation $\mathbf{x}_{T}$ corresponding to $\mathbf{x}$ by tracing the PF-ODE trajectory. In this work, we employ the PF-ODE derived from the DDIM framework (Song, Meng, and Ermon 2021), as shown in Equation 5. By applying an Euler approximation to Equation 5 and integrating forward, we obtain the following discrete update equation for latent inversion:
$\mathbf{x}_{t+1}$ $=\sqrt{\alpha_{t+1}}f_{\theta}(\mathbf{x}_{t})+\sqrt{1-\alpha_{t+1}}\,\epsilon_{\theta}^{(t)}\bigl{(}\mathbf{x}_{t}\bigr{)}.$ (6)
However, directly deriving $\mathbf{x}_{T}$ with Equation 6 involves iterative computation of $\epsilon_{\theta}^{(t)}$ for $T$ times, leading to slower inference. Here, we propose to only use the partial process of the original diffusion process in Equation 1. Given the discrete timestep set $\{1,2,\ldots,T\}$, we prepare a monotonically increasing subset $\bm{\tau}_{S}=[\tau_{1},\tau_{2},\ldots,\tau_{S}=T]\subset\{1,2,\ldots,T\}$, where $S\in\{1,2,\ldots,T\}$ is the number of the partial process. The inversion modeling can thus be accelerated by simply noising the test image with fewer timesteps stored in $\bm{\tau}_{S}$, which is given by
$\mathbf{x}_{\tau_{i+1}}$ $=\sqrt{\alpha_{\tau_{i+1}}}f_{\theta}(\mathbf{x}_{\tau_{i}})+\sqrt{1-\alpha_{\tau_{i+1}}}\,\epsilon_{\theta}^{(\tau_{i})}\bigl{(}\mathbf{x}_{\tau_{i}}\bigr{)}.$ (7)
Intuitively, Equation 7 leads to a considerable speed-up for the inference stage when $T\gg S$. Nevertheless, it also inevitably sacrifices approximation precision as the step interval for the Euler integration is extended. Recall here that our goal is to resolve AD rather than pursue accurate pixel-wise reconstruction. As such, we argue that high AD performance does not necessarily require reconstruction precision. It should be noticed that, even with limited inversion steps, the inversion mechanism in projecting an anomaly pixel can nonetheless yield sufficiently low typicality. In this regard, despite the poorly reconstructed samples, an anomaly pixel can still be expectedly mapped to a low-density region within the noise prior to achieve high detection performance. This insight suggests that our inversion-based AD method is free from the accuracy-efficiency trade-off in conventional reconstruction-based methods. We next need to find an ideal space that best suits our diffusion-based AD modeling.
Algorithm 1 Inference with InvAD
[…]
2:Anomaly score $s$, Anomaly map $A\in\mathbb{R}^{H\times W}$
3:/* DDIM Inversion */
4:Initialize $\mathbf{z}\leftarrow g_{\phi}(\mathbf{x}_{0})\in\mathbb{R}^{C\times h\times w}$
5:for $i=0$ to S1S-1 do
6:  Update $\mathbf{z}$ with Equation  7
7:  if i=S1i=S-1 then
8:   break
9:  end if
10:end for
11:/* Anomaly scoring */
12:Initialize $A\leftarrow\mathbf{0}\in\mathbb{R}^{h\times w}$
13:for $u\in\{0,\ldots,h-1\},v\in\{0,\ldots,w-1\}$ do
14:  Update $A[u,v]$ with Equation 8
15:end for
16:Upsample $A\in\mathbb{R}^{h\times w}$ to $\mathbb{R}^{H\times W}$
[…]
18:return $(s,A)$
Diffusion Modeling in Feature Space
Current image AD methods (Batzner, Heckler, and König 2023; Roth et al. 2021; Um et al. 2023; Mousakhan, Brox, and Tayyub 2023) heavily rely on pre-trained networks, referred to as backbone models, that are trained on large-scale datasets such as ImageNet (Deng et al. 2009). In diffusion-based AD, these backbones are used to compare the original test image with its reconstructed counterpart in feature space (Mousakhan, Brox, and Tayyub 2023; Lu et al. 2023a; He et al. 2024). Because backbone features exhibit invariance to low-level variations (e.g., noise or lighting conditions), they can significantly enhance AD performance.
However, in our inversion-based approach, the inferred latent $\mathbf{x}_{T}$ follows a standard Gaussian distribution, not the data distribution. To nonetheless utilize backbone models, we design our method to follow the latent-diffusion manner by extracting the feature $\mathbf{z}$ via $\mathbf{z}=g_{\phi}(\mathbf{x})\in\mathbb{R}^{C\times h\times w}$, where $g_{\phi}(\cdot)$ denotes the backbone model, and $C,(h,w)$ mean the feature map dimension and resolutions, respectively. This feature-space diffusion modeling enables our inference stage to refer to high-level semantics, rather than low-level pixel patterns. Also, such a latent-diffusion-based framework leads to further inference efficiency.
Table 2: Quantitative results on different AD datasets under the multi-class setting. The best and the second-best results are highlighted in bold and underlined, respectively. For each row, the mAD averages the evaluation over all image- and pixel-level metrics (i.e., the 3rd to the 10th column). FPS measures the image-level detection efficiency.
Dataset Method Image-level Pixel-level mAD FPS
AU-ROC AP F1_max AU-ROC AP F1_max AU-PRO
MVTecAD RD4AD [CVPR’22] 94.6 96.5 95.2 96.1 48.6 53.8 91.1 82.3 4.8
UniAD [NeurIPS’22] 96.5 98.8 96.2 96.8 43.4 49.5 90.7 81.7 5.3
SimpleNet [CVPR’23] 95.3 98.4 95.8 96.9 45.9 49.7 86.5 81.2 N/A
DeSTSeg [CVPR’23] 89.2 95.5 91.6 93.1 54.3 50.9 64.8 77.1 N/A
DiAD [AAAI’24] 97.2 99.0 96.5 96.8 52.6 55.5 90.7 84.0 0.1
HVQ-Trans [NeurIPS’23] 98.0 99.5 97.5 97.3 48.2 53.3 91.4 83.6 5.6
OmiAD [ICML’25] 98.8 99.7 98.5 97.7 52.6 56.7 93.2 85.3 39.4
InvAD (Ours) 99.0 99.6 98.5 97.5 46.5 52.3 92.7 83.7 88.1
VisA RD4AD [CVPR’22] 92.4 92.4 89.6 98.1 38.0 42.6 91.8 77.8 4.9
UniAD [NeurIPS’22] 88.8 90.8 85.8 98.3 33.7 39.0 85.5 74.6 4.6
SimpleNet [CVPR’23] 87.2 87.0 81.8 96.8 34.7 37.8 81.4 72.4 N/A
DeSTSeg [CVPR’23] 88.9 89.0 85.2 96.1 39.6 43.4 67.4 72.8 N/A
DiAD [AAAI’24] 86.8 88.3 85.1 96.0 26.1 33.0 75.2 70.1 0.1
HVQ-Trans [NeurIPS’23] 93.2 92.8 87.6 98.7 35.0 39.6 86.3 76.2 5.0
OmiAD [ICML’25] 95.3 96.0 91.2 98.9 40.4 44.1 89.2 79.3 35.3
InvAD (Ours) 96.9 97.2 93.7 99.1 39.2 43.1 92.9 80.3 74.1
MPDD RD4AD [CVPR’22] 84.1 83.2 84.1 98.1 35.2 38.7 93.4 73.8 4.7
UniAD [NeurIPS’22] 82.2 87.1 85.1 95.1 18.9 25.0 81.9 67.9 5.8
SimpleNet [CVPR’23] 90.6 94.1 89.7 97.1 33.6 35.7 90.0 75.8 N/A
DeSTSeg [CVPR’23] 93.0 95.1 90.6 94.1 33.2 37.6 59.8 71.9 N/A
DiAD [AAAI’24] 74.6 82.1 82.5 93.0 15.9 21.2 78.4 64.0 0.1
HVQ-Trans [NeurIPS’23] 86.5 88.1 85.8 96.7 27.6 31.4 86.9 71.9 6.2
OmiAD [ICML’25] 93.7 95.5 90.9 98.6 37.6 42.3 94.0 78.9 49.8
InvAD (Ours) 96.5 96.5 94.4 98.3 40.1 43.9 91.2 80.1 120
Anomaly Score Calculation
Once we obtain the latent representation $\mathbf{z}_{T}\in\mathbb{R}^{C\times h\times w}$ for features of a given test sample $\mathbf{x}$, the most straightforward way to compute the anomaly score is via the log-likelihood under the latent Gaussian distribution. However, it often triggers the known reverse-scoring problem and fails to capture small anomalies (See supplementary materials). Given that anomalies are inherently sparse in images, we additionally propose a norm-based image-level anomaly scoring approach. Specifically, we compute the Euclidean norm over all channels $C$ for each pixel $(i,j)$ in $\mathbf{z}_{T}$, which is written as
$\mathbf{z}_{T}^{\text{normed }}[i,j]=\|\mathbf{z}_{T}[:,i,j]\|_{2}\quad.$ (8)
Different from (He et al. 2024), for the image-level anomaly score, we compute the absolute difference between the maximum and minimum values of $\mathbf{z}_{T}^{\text{normed}}$. This is motivated by the observation that, even if an image is anomalous, anomalies for images are often locally and sparsely distributed within small regions, and the majority of image regions remain normal. In essence, our scheme for image-level anomaly scoring examines the difference between the highest- and the lowest-norm pixel to mitigate the influence of potential outlier pixels, thereby alleviating the reverse-scoring issue. The pixel-level anomaly score map $M_{\text{px}}$ is simply produced by interpolating $\mathbf{z}_{T}^{\text{normed}}\in\mathbb{R}^{h\times w}$ to match the desired output resolution $(H,W)$.
Our proposed InvAD paradigm is summarized in Algorithm 1. It is efficient yet effective, and can be applied in a plug-and-play manner for existing diffusion-based AD models, as it is devised only for the inference stage.
Table 3: Quantitative results per category on MVTecAD under the single-class setting, where we report (image-level AU-ROC, pixel-level AU-ROC) and FPS. The best results per category are highlighted in bold.
Category DiffAD GLAD TransFusion DAD DDAD ScoreDD DeCoDiff InvAD (Ours)
Texture carpet (98.3, 98.1) (99.0, 98.5) (99.2, 98.9) (99.3, 99.0) (99.3, 98.7) (96.9, 98.9) (85.2, 96.3) (100, 98.0)
grid (100, 99.7) (100, 99.6) (100, 99.1) (100, 99.7) (100, 99.4) (100, 99.7) (100, 99.5) (100, 97.4)
leather (100, 99.1) (100, 99.8) (100, 99.5) (100, 99.8) (100, 99.4) (99.6, 99.3) (97.5, 98.8) (100, 98.0)
tile (100, 99.4) (100, 98.7) (99.8, 92.1) (99.9, 99.7) (100, 98.2) (98.6, 94.4) (97.4, 98.2) (100, 94.4)
wood (100, 96.7) (99.4, 98.4) (99.4, 94.5) (99.9, 96.7) (100, 95.0) (98.8, 95.1) (94.6, 97.3) (99.3, 92.3)
Object bottle (100, 98.8) (100, 98.9) (100, 97.7) (99.7, 99.1) (100, 98.7) (100, 97.9) (100, 98.7) (100, 97.7)
cable (94.6, 96.8) (99.9, 98.1) (97.9, 95.6) (99.6, 98.1) (99.4, 98.1) (96.8, 97.5) (99.2, 98.3) (99.7, 98.7)
capsule (97.5, 98.2) (99.5, 98.5) (98.5, 97.5) (98.9, 98.3) (99.4, 95.7) (96.1, 98.6) (91.0, 97.2) (99.4, 98.8)
hazelnut (100, 99.4) (100, 99.5) (100, 97.3) (100, 99.7) (100, 98.4) (99.9, 99.2) (99.5, 97.8) (99.9, 98.1)
metal nut (99.5, 99.1) (100, 98.8) (100, 96.8) (100, 99.7) (100, 99.0) (97.2, 97.9) (97.9, 97.9) (99.7, 95.4)
pill (97.7, 97.7) (98.1, 97.9) (98.3, 92.5) (99.0, 99.3) (100, 99.1) (95.3, 96.0) (96.4, 98.9) (98.7, 98.1)
screw (97.2, 99.0) (96.9, 99.1) (97.2, 99.0) (99.4, 99.0) (99.0, 99.3) (99.6, 99.6) (86.7, 98.3) (98.1, 99.6)
toothbrush (100, 99.2) (100, 99.4) (100, 98.6) (100, 98.8) (100, 98.7) (99.8, 98.3) (91.0, 97.2) (92.8, 97.8)
transistor (96.1, 93.7) (98.3, 96.2) (98.3, 93.1) (99.8, 92.9) (100, 95.3) (95.4, 95.2) (99.8, 97.9) (100, 97.9)
zipper (100, 99.0) (98.5, 97.9) (100, 97.6) (100, 99.2) (100, 98.2) (99.8, 99.3) (99.2, 99.5) (99.2, 97.4)
Avg. (98.7, 98.3) (99.3, 98.6) (99.2, 96.7) (99.7, 98.7) (99.8, 98.1) (98.2, 97.8) (95.7, 98.1) (99.1, 97.3)
FPS N/A 0.2 1.6 9.1 0.8 N/A 17 88.1
Experiments
Datasets
We use three widely-used unsupervised AD benchmarks—MVTecAD (Bergmann et al. 2019), VisA (Zou et al. 2022), and MPDD (Jezek et al. 2021)—all designed for measuring real-world AD performance on RGB industrial images. Since each benchmark provides both image- and pixel-level annotations, we can evaluate performance in terms of overall AD (image-level) as well as precise localization (pixel-level). We provide a detailed description of these datasets in the supplementary materials.
Evaluation Metrics
Following (Feng et al. 2025), we employ Area Under the Receiver Operating Characteristic Curve (AU-ROC), Average Precision (AP), F1-score max (F1_max) for both image- and pixel-level AD evaluation. We also adopt the Area Under the Per-Region-Overlap (AU-PRO) (Bergmann et al. 2019) metric to examine anomaly localization performance at the pixel level. Furthermore, we apply Frames Per Second (FPS) to evaluate the efficiency of inference.
Implementations
To assess the effectiveness of the proposed method, we begin by training a Diffusion Transformer (DiT)-based diffusion model (Peebles and Xie 2023) from scratch on each dataset with Equation 3, where we set $\gamma=\mathbf{1}$ as DDPM (Ho, Jain, and Abbeel 2020). To ensure fair comparisons, we employ the pre-trained EfficientNet-B4 (Tan and Le 2019) model on ImageNet-1k (Deng et al. 2009) to encode the images into feature space as OmiAD (Feng et al. 2025) and HVQ-Trans (Lu et al. 2023b). We train the diffusion model for 300 epochs across all datasets using the AdamW optimizer (Loshchilov and Hutter 2017) with a cosine learning rate schedule. The total diffusion step $T$ is set to 1000 during training using a linear noise schedule. For all experiments, we uniformly set $S=3$ to skip intermediate steps within the original $T$ steps, and create the subset $\tau_{3}=[333,666,999]$ to acclerate inversion compuatation. We conduct all throughput measurements on an NVIDIA RTX 4090 (24GB) with batch size set to 64. Note that the fluctuation in throughput measurements on different datasets is because some small test sets contain test samples less than the batch size (<64<64), where we also use a smaller batch size setting. See supplementary materials for more details.
Table 4: Comparison of reconstruction-based approach (Recon.) against our inversion-based AD paradigm, under different total diffusion steps $S$ and perturbation ratios $r$, in multi-class MVTecAD with AU-ROC.
Ratio $r$ Total Diffusion Steps $S$
3 5 10 50 100 1000
Recon.
10 % N/A N/A 66.6 97.7 97.9 98.0
20 % N/A 64.7 97.9 91.3 98.2 98.2
40 % 64.9 75.0 89.4 98.0 98.2 98.2
60 % 64.9 68.4 87.1 97.3 97.8 98.1
80 % 67.8 67.8 74.1 91.8 92.1 93.9
Inversion (Ours) 99.0 98.9 98.4 96.0 95.7 95.4
Comparison with SOTA AD Methods
We compare the quantitative results against prior multi-class AD competitors, including (i) Transformer-based methods (HVQ-Trans (Lu et al. 2023b), UniAD (You et al. 2022)), (ii) reconstruction-based methods (RD4AD (Deng and Li 2022), DeSTSeg (Zhang et al. 2023c)), (iii) diffusion-based methods (DiAD (He et al. 2024), OmiAD (Feng et al. 2025)). The results are summarized in Table 2. It can be seen that our method generally yields the best image-level and highly competitive pixel-level detection performance on three datasets. Notably, our method consistently outperforms all the compared methods regarding inference efficiency by a large margin (i.e., FPS: 88.18 vs. 39.37 on MVTecAD). This can be attributed to the property of our proposed inversion-based AD paradigm, which enables high detection accuracy with fewer NFEs for acceleration. By contrast, conventional diffusion-based methods, such as DiAD (He et al. 2024), require VAE decoding for reconstruction in the pixel space, thus resulting in slower inference. It is noteworthy that our method is also faster than OmiAD whose NFE is 1, which is primarily due to the heavy diffusion backbone and its modeling strategy in high-resolution feature space.
Table 5: Generalizability evaluation of our method when incorporated into other diffusion-based AD approaches on MVTecAD and VisA with AU-ROC and FPS.
Method Image AU-ROC [%] FPS
MVTecAD VisA MVTecAD
DeCoDiff [CVPR’25] 99.3 96.1 17
DeCoDiff + InvAD 97.0 (-2.0) 93.2 (-2.8) 46 (+29)
MDM [ICML’25] 91.9 80.5 2.2
MDM + InvAD 98.2 (+6.3) 94.8 (+14.3) 63 (+60.8)
In addition, we compare our method against other prior single-class diffusion-based AD methods in Table 3, including DiffAD (Zhang et al. 2023b), GLAD (Yao et al. 2024), TransFusion (Fučka, Zavrtanik, and Skočaj 2025), DAD (Zhang et al. 2023a), DDAD (Mousakhan, Brox, and Tayyub 2023), ScoreDD (Teng et al. 2023), DeCoDiff (Beizaee et al. 2025). Similar to the multi-class scenario, our method consistently improves efficiency with competitive accuracy.
Is Inversion Effective?
Our approach builds on the inversion modeling to realize reconstruction-free AD. To investigate the validity of this design, we vary two key parameters in reconstruction-based methods, total diffusion steps $S$ and perturbation timesteps $t$, to gain a deeper insight into the effect. In particular, instead of inspecting $t$ directly, we study the ratio of $r=t/S$ as it better captures the perturbation strength across different $S$. The results are summarized in Table 4, where we only modify the inference stage to examine the accuracy (i.e., image-level AU-ROC) change. It can be confirmed that for reconstruction-based methods, the parameter tuning efforts can be highly burdensome to yield ideal performance. While a larger $S$ tends to perform better, it induces computationally expensive denoising iterations.
By contrast, our inversion-based method can achieve the best performance even with an extremely small inversion step setting (e.g., $S=3,5$ in the bottom row of Table 4), with no reliance on perturbation timesteps as it is not involved in our modeling. This further evidences the superiority of our InvAD for the task of image AD.
Generalizability Evaluation
As our InvAD is designed for the inference stage, it can be incorporated into existing diffusion-based AD models in a plug-and-play manner. To evaluate the effectiveness, we incorporate InvAD into the inference stage for DeCoDiff (Beizaee et al. 2025) and Masked Diffusion Model (MDM) (Feng et al. 2025), and report the image-level AU-ROC and the FPS metrics on both datasets. It can be seen from Table 5 that InvAD leads to FPS gain for both methods with only a minor sacrifice in accuracy for DeCoDiff. This suggests that InvAD is model-agnostic and can be inserted as a plug-and-play module into different diffusion-based AD techniques.
Also, it is worth mentioning that for MDM, our method also leads to an evident performance improvement regarding AU-ROC. We assume that this is because, different from DeCoDiff that predicts the direction of deviation, MDM learns to optimize the noise residual during training, which suits better for our method.
Table 6: Design ablation of feature space encoding modules on MVTecAD and VisA with AU-ROC and FPS.
Model Image AU-ROC [%] # Params FPS
MVTecAD VisA MVTecAD
VAE 87.1 86.9 1289M 52.6
ResNet-18 88.7 82.4 1239M 68.4
DINO 96.1 91.0 1312M 85.3
EfficientNet 99.0 96.9 1243M 88.1
Table 7: Design ablation of diffusion architectures on MVTecAD and VisA with AU-ROC and FPS.
Model Image AU-ROC [%] # Params FPS
MVTecAD VisA MVTecAD
MLP 97.2 91.1 351M 241
UNet 98.0 88.4 1433M 288
DiT-base 94.4 88.4 155M 240
DiT-gigant 99.0 96.9 1223M 88
Backbone Design Ablation
Our inversion-based method is considered in the pre-trained feature space rather than pixel space. In general, the selection of feature extractors heavily impacts the performance of AD methods. In Table 6, we evaluate the performance with different backbones including VAE (Rombach et al. 2022), ResNet-18 (He et al. 2016), DINO (Caron et al. 2021), and EfficientNet (Tan and Le 2019). It can be seen that stronger feature extractors (i.e., DINO-base and EfficientNet-B4) lead to better performance than those with CNN-based backbones (i.e., VAE and ResNet-18). Also, we study the influence of the design of diffusion modeling backbones on AD performance in Table 7. In particular, we compare four types of backbones: MLP, UNet, DiT-base, and DiT-gigant (Peebles and Xie 2023). In contrast to the feature encoding modules, we observe that the performance is generally similar, even with the simple MLP implementation. This suggests that our method is robust to diffusion architectures, but also relies on the quality of embeddings derived with feature extractors, as prior methods (Feng et al. 2025; Roth et al. 2021). See supplementary materials for a detailed description of backbones.
Figure 2: Visualization of anomaly localization on MVTecAD (1st row), VisA (2nd row), and MPDD (3rd row). GT displays the ground-truth anomaly map.
Visualization Analysis
We qualitatively evaluate the pixel-level AD performance of our inversion-based method against OmiAD (Feng et al. 2025), in Figure 2. Similar to the quantitative evaluation in Table 2, our method generally achieves comparable anomaly localization capability with OmiAD, which again, demonstrates that our method can well improve efficiency yet remain high accuracy.
Conclusion
We propose a simple yet effective diffusion-based AD paradigm, “detection via noising in latent space”, to address the fundamental challenges within the conventional “detection via denoising in RGB space”. Our method, InvAD, follows a reconstruction-free manner to directly infer the final latent state with the DDIM inversion mechanism. To pursue efficiency, we simply enforce very few inversion steps to noise a test sample, yet manage to maintain high detection accuracy as the typicality can still be accurately scored. Due to the reconstruction-free design, InvAD is inherently tuning-free for the perturbation timestep. We experimentally demonstrate that our method achieves state-of-the-art detection performance, especially regarding inference efficiency.
While our method can achieve satisfactory performance, it still needs more than one NFE. To further improve efficiency (i.e., only require 1 NFE), we will consider introducing diffusion distillation to compactly compress the progressive inversion process into only one single step.
Appendix A A. Dataset Details
We utilize three publicly available anomaly detection (AD) benchmarks for evaluation: MVTecAD (Bergmann et al. 2019), VisA (Zou et al. 2022), MPDD (Jezek et al. 2021). Table 8 summarizes the key statistics of the datasets used in our experiments, including the number of classes, domain, image resolution, and sample counts for training and testing. All of these datasets follow an unsupervised AD setting. In other words, the training set only consists of defect-free normal images, and the test set consists of both normal and anomalous images.
Appendix B B. Implementation Details
Data Preparation
We resize all training images to $\times256$ pixels and do not apply any data augmentation. Image normalization is performed using the ImageNet dataset’s channel-wise mean ([0.485, 0.456, 0.406]) and standard deviation ([0.229, 0.224, 0.225]).
Model Setup
We adopt the original DiT architecture (Peebles and Xie 2023), but remove all class‐conditioning so that our method does not depend on class labels. This design allows DiT to be applied in scenarios where ground‐truth annotations are unavailable or impractical (e.g., inspection of diverse products on a manufacturing line). We set the number of transformer blocks and embedding dimensions to 16 and 2048 for DiT-gigant, and to 8 and 1024 for DiT-base, respectively. For the UNet‐based diffusion models evaluated in Table 7, we follow the architectural specifications of the original DDPM implementation (Ho, Jain, and Abbeel 2020), consistent with prior work (Beizaee et al. 2025; He et al. 2024). For the MLP‐based diffusion models, we stack multiple MLP blocks and apply AdaLN-Zero conditioning (Peebles and Xie 2023), as in (Li et al. 2025). We summarize all major hyperparameters for diffusion modeling in Table 9.
Table 8: Dataset statistics used in our experiments. ‘N’ and ‘A’ denote the numbers of normal and anomalous samples, respectively.
MVTecAD VisA MPDD
Classes 15 12 6
Domain Industrial Industrial Metal
Resolution $\times1024$ $\times256$ $\times1024$
Train (N) 2,629 8,659 888
Test (N) 467 962 176
Test (A) 1,258 1,200 282
Table 9: Hyperparameters of diffusion modeling.
Parameter Value
Diffusion Settings
Training diffusion steps ($T$) 1000
Inference diffusion steps ($S$) 3
Noise scheduler Linear, $\beta_{1}=1\times 10^{-4}$, $\beta_{T}=0.02$
Parameterization $\epsilon$-prediction
Noise scale ($\sigma$) Fixed
Training Settings
Batch size 8
Optimizer AdamW
Gradient clipping threshold 1.0
Initial learning rate $1\times 10^{-6}$
Peak learning rate $5\times 10^{-5}$
Final learning rate $5\times 10^{-6}$
Learning rate scheduler Warmup + cosine decay
Warmup epochs 40
Total number of epochs 300
Weight decay 0
Evaluation Pipeline
We evaluate AD performance at both image- and pixel-levels using the following metrics:
  • Area Under the Receiver Operating Characteristic Curve (AU-ROC) provides an intuition on how well the model takes a trade-off between recall and precision in a threshold-independent manner.
  • Average Precision (AP) conducts a more practical evaluation than AU-ROC, where the test set contains far fewer anomalous samples compared to normal ones.
  • Maximum F1-score ($\bm{\mathrm{F1}_{\max}}$) measures the F1 score on the best-performed threshold, which reflects actual performance on the real problem.
To assess pixel-level anomaly localization, we additionally employ:
  • Area Under the Per-Region-Overlap Curve (AU-PRO) calculates the IoU between prediction and anomalous region, which is more robust to the small anomalies (Bergmann et al. 2019).
Experimental Environment
Our implementations are mostly based on Pytorch-2.7.1 and Python-3.10. We show the machine specification for the training and evaluation in our experiments in Table 10.
Table 10: Experimental environments for training and evaluation.
Component Training Evaluation
CPU Xeon 8468 ×2 i9-14900KF
GPU NVIDIA H200 ×8 RTX 4090 (24GB)
RAM 1.0 TiB ($\approx$1024GB) 31 GB
OS RHEL 9.4 (Plow) Ubuntu 24.04.2 LTS
To ensure reproducibility, we fixed random seeds for all libraries (NumPy, PyTorch, etc.) and disabled nondeterministic GPU operations if possible.
Appendix C C. Additional Quantitative Results on Industrial Anomaly Detection
To further investigate the effectiveness of our proposed framework, we evaluate in Table 11 the detection performance on VisA (Zou et al. 2022) in a single-class setting. In addition to the state-of-the-art performance on the multi-class AD setting explained in our manuscript (Table 3), our InvAD also shows competitive single-class detection accuracy with a remarkable efficiency gain.
Table 11: Quantitative results per category on VisA under the single-class setting, where we report (image-level AU-ROC, pixel-level AU-ROC) and FPS. The best results per category are highlighted in bold.
Category GLAD TransFusion DAD DDAD InvAD (Ours)
candle (99.9, 94.8) (98.3, N/A) (98.7, N/A) (99.9, 98.7) (98.0, 99.4)
capsules (99.1, 99.6) (99.6, N/A) (97.9, N/A) (100, 99.5) (92.1, 99.7)
cashew (98.4, 97.0) (93.7, N/A) (96.5, N/A) (94.5, 97.4) (95.6, 98.9)
chewinggum (99.6, 99.1) (99.6, N/A) (99.9, N/A) (98.1, 96.5) (99.7, 98.4)
fryum (99.4, 96.9) (98.3, N/A) (98.3, N/A) (99.0, 96.9) (97.5, 96.9)
macaroni1 (99.9, 99.8) (99.4, N/A) (99.5, N/A) (99.2, 98.7) (97.8, 99.7)
macaroni2 (98.9, 99.8) (96.5, N/A) (99.0, N/A) (99.2, 98.2) (86.9, 99.4)
pcb1 (99.6, 99.6) (98.9, N/A) (99.2, N/A) (100, 93.4) (99.5, 99.8)
pcb2 (100, 98.6) (99.7, N/A) (99.1, N/A) (99.7, 97.4) (99.5, 99.1)
pcb3 (99.9, 98.9) (99.2, N/A) (98.6, N/A) (97.2, 96.3) (98.6, 99.2)
pcb4 (99.9, 99.5) (99.6, N/A) (98.9, N/A) (100, 98.5) (99.8, 98.3)
pipe fryum (98.9, 99.4) (99.6, N/A) (99.8, N/A) (100, 99.5) (99.1, 99.2)
Avg. (99.5, 98.6) (98.5, 88.8) (98.8, 98.9) (98.9, 97.6) (97.0, 99.0)
FPS 0.2 1.6 9.1 0.8 88.1
Appendix D D. Generalizability on Other Domains
While our InvAD achieves state-of-the-art performance on industrial image AD benchmarks (e.g., MVTecAD, VisA, MPDD), it is still unclear whether it generalizes to other AD tasks like medical imaging. To address this, we evaluate InvAD on a representative medical imaging benchmark, BMAD (Bao et al. 2024), which is a comprehensive AD benchmark including six reorganized datasets supporting medical AD domains, such as brain MRI, liver CT, and retinal OCT. The BMAD consists of 52,742 normal training images, 22,632 anomalous test images, and 4,540 normal test images. We report image-level AU-ROC (det.), and pixel-level AU-ROC (loc.), if the ground-truth anomaly maps are provided. For the comparison with other AD methods, we directly report the results in BMAD (Bao et al. 2024).
We compare quantitative results per dataset on BMAD under the single-class setting in LABEL:tab:sup_bmad. In summary, our proposed InvAD achieves state-of-the-art performance on the BMAD benchmark with huge improvements in inference efficiency. These results further support the effectiveness of our proposed framework, not only within the domain of industrial inspection but also in other fields, such as medical diagnosis.
Table 12: Quantitative results on BMAD under the single-class setting, where we report image-level AU-ROC (det.), pixel-level AU-ROC (loc.), and FPS. The best and the second-best results are highlighted in bold and underlined, respectively. For each row, the mAD averages the evaluation over all image- and pixel-level metrics. FPS measures the image-level detection efficiency on the RESC dataset.
Methods BraTS2021 BTCV + LiTS RESC OCT2017 RSNA Camelyon16 mAD FPS
det. loc. det. loc. loc. loc. det. det. det.
PaDiM [ICPR’20] 79.0 94.4 50.8 90.9 75.9 91.4 91.8 77.5 67.3 79.9 20.4
CFlow [CVPR’20] 74.8 93.8 50.8 92.4 75.0 93.8 85.4 71.5 55.7 77.0 15.1
RD4AD [CVPR’22] 89.5 96.5 60.4 96.0 87.8 96.2 97.3 67.6 66.8 84.2 22.6
PatchCore [CVPR’22] 91.7 97.0 60.3 96.4 91.6 96.5 98.6 76.1 69.3 86.4 20.1
SimpleNet [CVPR’23] 82.5 94.8 72.3 97.5 76.2 77.1 94.7 69.1 62.4 80.7 10.8
InvAD (Ours) 90.3 97.8 60.6 97.8 93.6 96.5 99.2 82.7 67.7 87.4 64.3
Appendix E E. Additional Background of Diffusion Models
Denoising Diffusion Probabilistic Models
Denoising diffusion probabilistic models (DDPM) (Ho, Jain, and Abbeel 2020) are a class of latent generative models which involve progressively collapsed $T$ random variables $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$. Let us assume we have i.i.d. data samples $\mathcal{X}=\{\mathbf{x}_{0}\mid\mathbf{x}_{0}\sim q_{0}(\mathbf{x})\}$ where $q_{0}(\mathbf{x})$ is an unknown data distribution. DDPM aims to approximate this data distribution by tracing the data collapsing process (forward process) in reverse order. In general, the forward process $q\bigl{(}\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\bigr{)}$ is represented as a Markov process with the Gaussian transition kernel:
$q\bigl{(}\mathbf{x}_{1:T}\mid\mathbf{x}_{0}\bigr{)}$ $:=\prod_{t=1}^{T}q\bigl{(}\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\bigr{)},$ (9)
$\text{where}\quad q\bigl{(}\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\bigr{)}$ $:=\mathcal{N}\!\Bigl{(}\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}\,\mathbf{x}_{t-1},\bigl{(}1-\frac{\alpha_{t}}{\alpha_{t-1}}\bigr{)}\mathbf{I}\Bigr{)}.$ (10)
The noise schedule is defined by the monotonically decreasing sequences $\{\alpha_{t}\}_{t=1}^{T}$ in the above equation. Notably, from the reproducibility of the Gaussian distribution, we have:
$q\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{0},\left(1-\alpha_{t}\right)\mathbf{I}\right).$ (11)
Since we can efficiently compute latent variables $\mathbf{x}_{t}$ given a clean sample $\mathbf{x}_{0}$, it provides a GPU-friendly diffusion model training. Also, we define the generative process $p_{\theta}\left(\mathbf{x}_{0:T}\right)$ as the Markov process with a learnable Gaussian transition kernel:
$p_{\theta}\left(\mathbf{x}_{0:T}\right):=p_{\theta}\left(\mathbf{x}_{T}\right)\prod_{t=1}^{T}p_{\theta}^{(t)}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right),$ (12)
$\text{where}\quad p_{\theta}^{(t)}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}):=\mathcal{N}\bigl{(}\mathbf{x}_{t-1};\bm{\mu}_{\theta}^{(t)}(\mathbf{x}_{t}),\sigma_{t}^{2}\mathbf{I}\bigr{)}.$ (13)
In DDPM, the posterior mean is modeled using a noise prediction neural network $\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})$:
$\bm{\mu}_{\theta}^{(t)}(\mathbf{x}_{t})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\bigl{(}\mathbf{x}_{t}-\frac{1-(\alpha_{t}/\alpha_{t-1})}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})\bigr{)}.$ (14)
The diffusion objective is expressed as:
$\mathcal{L}_{\gamma}\left(\epsilon_{\theta}\right)\coloneqq\sum_{t=1}^{T}\gamma_{t}\mathbb{E}_{\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right),\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left\|\epsilon_{\theta}^{(t)}(\mathbf{x}_{t})-\epsilon_{t}\right\|_{2}^{2}\right],$ (15)
where $\mathbf{x}_{t}$ is efficiently sampled by Equation 11, and $\mathbf{\gamma}\coloneqq[\gamma_{1},\ldots,\gamma_{T}]^{T}$ denotes the weighting vector.
Denoising Diffusion Implicit Models
The forward process in Equation 9 is a Markov process where the current state $\mathbf{x}_{t}$ solely depends on the previous state $\mathbf{x}_{t-1}$. Denoising diffusion implicit models (DDIM) (Song, Meng, and Ermon 2021) is a non-Markov diffusion model where each forward process is conditioned on both $\mathbf{x}_{0}$ and $\mathbf{x}_{t-1}$. Specifically, the forward posterior is defined as:
$q_{\sigma}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\mathbf{x}_{0}\right)\\ =\mathcal{N}\Biggl{(}\sqrt{\alpha_{t-1}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\frac{\mathbf{x}_{t}-\sqrt{\alpha_{t}}\mathbf{x}_{0}}{\sqrt{1-\alpha_{t}}},\sigma_{t}^{2}\bm{I}\Biggr{)},$ (16)
where $\{\sigma_{t}\}_{t=1}^{T}$ controls the uncertainly of the diffusion trajectories. Since this posterior is designed to match $q(\mathbf{x}_{t}\mid\mathbf{x}_{0})$ as DDPM, the generative process can be trained with the same objective described in Equation 15, for any $\{\sigma_{t}\}_{t=1}^{T}$. When $\sigma_{t}\rightarrow 0$, the generative process becomes deterministic and can be represented as:
$\mathbf{x}_{t-1}=\sqrt{\alpha_{t-1}}\bm{f}_{\theta}\left(\mathbf{x}_{t},t\right)+\sqrt{1-\alpha_{t-1}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t}),$ (17)
where $\bm{f}_{\theta}(\mathbf{x}_{t},t)=(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}^{(t)}(\mathbf{x}_{t}))/\sqrt{\alpha_{t}}$. For more detailed proof and derivation of DDIM, see (Song, Meng, and Ermon 2021).
Derivation of DDIM Inversion
From Equation 17, we have:
$\mathbf{y}_{t-1}-\mathbf{y}_{t}=(p_{t-1}-p_{t})\epsilon_{\theta}^{(t)}(\mathbf{x}_{t}),$ (18)
where $\mathbf{y}_{t}\coloneqq\mathbf{x}_{t}/\sqrt{\alpha_{t}}$ and $p_{t}\coloneqq\sqrt{1/{\alpha_{t}}-1}$. In the limit of $T\rightarrow\infty$, the above differential equation converges to a continuous-time ODE:
$\mathrm{d}\mathbf{y}_{t}=\epsilon_{\theta}^{(t)}\,\mathrm{d}p_{t}.$ (19)
Based on the ODE formulation of DDIM, we can derive Equation 17 by applying Euler integration to Equation 19 with the step size $\Delta p_{t}^{+}:=p_{t-1}-p_{t}(>0)$. Similarly, by applying the forward Euler method to the Equation 19 within the range [t1,t][t-1,t], we have:
$\mathbf{y}_{t}=\mathbf{y}_{t-1}+\Delta p_{t}^{-}\epsilon_{\theta}^{(t-1)}(\mathbf{x}_{t-1}),$ (20)
where $\Delta p_{t}^{-}:=p_{t}-p_{t-1}$. Next, replacing $\mathbf{y}_{t}$ with $\mathbf{x}_{t}/\sqrt{\alpha}_{t}$ in Equation 20 yields:
$\mathbf{x}_{t}=\sqrt{\alpha_{t}}\bigl{(}\frac{\mathbf{x}_{t-1}}{\sqrt{\alpha_{t-1}}}+\Delta p_{t}^{-}\epsilon_{\theta}^{(t-1)}(\mathbf{x}_{t-1})\bigr{)}.$ (21)
It can be easily noticed that Equation Equation 21 is essentially equivalent to Equation (6). A similar discussion applies to the subset of timesteps $\bm{\tau}_{S}=[\tau_{1},\tau_{2},\ldots,\tau_{S}=T]\subset\{1,2,\ldots,T\}$ by considering arbitray time interval $\tau_{s},\tau_{s+1}]$.
Appendix F F. Ablation on Scoring Scheme
As explained in the paragraph “Anomaly Score Calculation” of our manuscript, our anomaly scoring scheme considers both negative-log-likelihood (NLL) in the noise prior and the spatial absolute difference of Euclidean norm (Abs) to alleviate the reverse-scoring problem (Nalisnick et al. 2019; Osada et al. 2023). To study the advantage of this strategy, we compare the performance in Table 13 with the conventional NLL-based scoring on MVTecAD. It can be seen that our scoring scheme (i.e., NLL + Abs) contributes to robustness to the diffusion step setting $S$, while the NLL or Abs solely do not achieve this. This demonstrates that our scoring scheme plays a key role in the proposed InvAD paradigm to realize reconstruction-free AD, where $S$ can be further set to be very small for efficiency.
Table 13: Effect of the scoring scheme, under different total diffusion steps $S$ in multi-class MVTecAD with image-lavel AU-ROC.
Total Diffusion Steps $S$ 3 5 10 50 100 1000
NLL 96.1 95.2 93.0 89.7 89.4 89.1
Abs 98.4 98.1 95.8 83.2 82.2 82.0
NLL + Abs (Ours) 99.0 98.9 98.4 96.0 95.7 95.4
Appendix G G. Additional Visualizations
We exhibit more visualization against qualitative AD results on MVTecAD and VisA, in Figure 3 and Figure 4.
Figure 3: Additional visualization of anomaly localization on MVTecAD.
Figure 4: Additional visualization of anomaly localization on VisA.